OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9 @Nanalew

OpenStudy (anonymous):

i think your in the wrong section but anyways, 7(5 + 9) = 7 · x + 7 · 9 simplify the ones in the bracket 7(14)=7x+7•9 then do multiplication 98=7x + 63 collect like terms 98-63=7x subtract 35=7x divide both sides by 7 to simplify 5=x therefore x=5

OpenStudy (anonymous):

lol, there you go ;p

OpenStudy (anonymous):

sorry...

OpenStudy (anonymous):

np :)